# HAODiff: Human-Aware One-Step Diffusion via Dual-Prompt Guidance

### v1

@@ Abstract
Human-centered images often suffer from severe generic degradation during transmission and are prone to human motion blur (HMB), making restoration challenging. Existing research lacks sufficient focus on these issues, as both problems often coexist in practice. To address this, we design a degradation pipeline that simulates the coexistence of HMB and generic noise, generating synthetic degraded data to train our proposed HAODiff, a human-aware one-step diffusion. Specifically, we propose a triple-branch dual-prompt guidance (DPG), which leverages high-quality images, residual noise (LQ minus HQ), and HMB segmentation masks as training targets. It produces a positive–negative prompt pair for classifier-free guidance (CFG) in a single diffusion step. The resulting adaptive dual prompts let HAODiff exploit CFG more effectively, boosting robustness against diverse degradations. For fair evaluation, we introduce MPII-Test, a benchmark rich in combined noise and HMB cases. Extensive experiments show that our HAODiff surpasses existing state-of-the-art (SOTA) methods in terms of both quantitative metrics and visual quality on synthetic and real-world datasets, including our introduced MPII-Test. Code is available at: https://github.com/gobunu/HAODiff.
@@ 1 Introduction
Human body restoration (HBR) focuses on recovering high-quality (HQ) images from low-quality (LQ) inputs that contain human images. When human subjects appear prominently in an image, it naturally attracts more viewer attention. However, real-world images frequently undergo degradation during capture or transmission, including human motion blur, noise, resolution loss, and JPEG compression artifacts. These distortions significantly hinder the recognition of human activities and limit the usefulness of such images in broader applications. As a result, many downstream tasks related to humans are negatively affected, such as 3D reconstruction wang20213Dreview ; zhang2025multigo , human pose estimation Samkari2023HPEreview ; zheng2025hipart , and human-object interaction detection wang2024hoidreview ; liu2025interacted .
To achieve better performance in practical scenarios, current blind image restoration (BIR) models typically rely on a large number of paired LQ and HQ images to accurately learn the complex mapping between LQ and HQ domains. Models in the HBR field follow the same principle. However, collecting a large number of real LQ-HQ image pairs is challenging. Real-world LQ images often undergo various unknown degradations, which may occur during transmission or even at the time of capture. Therefore, BIR models usually adopt a degradation pipeline to simulate real-world LQ conditions. This strategy allows models to fully leverage large-scale high-quality image restoration datasets, such as LSDIR Li2023LSDIR , FFHQ karras2019ffhq , and PERSONA gong2025osdhuman . In the context of HBR, existing approaches gong2025osdhuman ; Zhang2024DiffBody generally use generic degradation types, including downsampling, compression, noise, and low-pass blur. A commonly adopted pipeline is from Real-ESRGAN wang2021real , which is originally developed for blind super-resolution tasks. However, this pipeline may fail to cover the diverse degradation types that frequently occur in human-centric images.
@@
Among the various degradation types in human-centric images, human motion blur is particularly prevalent, yet it remains significantly underrepresented in existing degradation pipelines. Motion blur can be broadly categorized into two types. The first category is global motion blur, typically caused by camera movement during exposure. The second is local motion blur, caused by the rapid movement of objects in the scene. In human images, local motion blur mainly arises from human movement, known as human motion blur (HMB). Extensive research is conducted on the removal of global motion blur. For example, DeblurGAN Kupyn2018DeblurGAN not only proposes an effective model for motion deblurring, but also provides a simulation pipeline that can generate realistic global motion blur. This enables training with large-scale synthetic data. In contrast, the restoration of local motion blur is typically based on real datasets, with ReLoBlur li2023real being a representative example that covers diverse object motion. However, these datasets often focus on general objects and pay less attention to human-specific motion blur. Moreover, HMB in real-world human images often co-occurs with other degradation types. Existing deblurring models are usually designed to handle a single degradation, which limits their applicability in more complex, realistic scenarios.
@@
Multiple factors, including training data and degradation pipelines, influence the quality of image restoration. Among them, the model architecture plays a critical role. In BIR tasks, models require strong generative capabilities to reconstruct damaged regions. Generative adversarial networks (GANs) goodfellow2014gan are considered one of the major starting points for modern image generation models. They have led to the development of many powerful restoration methods zhang2021bsrgan ; Kupyn2018DeblurGAN ; wang2021real . However, GANs often suffer from training instability and difficulties in controlling the generation process. Diffusion models and latent diffusion models (LDMs) Rombach2022LDM have further improved the framework for both generation and restoration. Some models lin2024diffbir ; yang2023pasd ; wu2024seesr ; yue2023resshift use multi-step diffusion processes to restore high-quality images from heavily degraded inputs. Recently, one-step diffusion models wu2024osediff ; wang2024osdface ; wang2024sinsr ; yue2024invsr also demonstrate strong BIR performance while significantly reducing resource consumption compared to multi-step diffusion. These models commonly extract object features from images and convert them into text or image embeddings. They are used as positive prompts to guide the model toward faithful restoration. To better leverage the guidance capabilities of text-to-image (T2I) foundation models, some methods 2024s3diff ; yu2024supir introduce negative prompts using classifier-free guidance (CFG) ho2021cfg . These prompts help steer the model away from undesired content. However, these methods adopt fixed negative prompts, which limit their guidance effectiveness.
@@
To address these limitations, we propose HAODiff, a novel one-step diffusion for human body restoration (HBR) that integrates CFG via dual-prompt guidance (DPG). Firstly, to compensate for the lack of human motion blur (HMB) in existing pipelines, we propose a new one. In the preprocessing stage, we apply a human body-part segmentation model to the HQ training images to obtain part-specific masks. These masks are then combined with a motion blur simulation module to synthesize HMB. Integrated into a generic two-stage pipeline, our approach allows the inclusion of HMB. Secondly, we design a triple-branch dual-prompt guidance, named DPG, which is based on the Swin Transformer liu2021Swin . During training with the degradation pipeline, one branch of DPG predicts the HQ image to produce the positive prompt. The other two are used to predict the residual noise (LQ minus HQ) and the human motion blur segmentation mask, both serving as sources for the negative prompt. Thirdly, we propose a human-aware one-step diffusion via dual-prompt guidance, named HAODiff. The model accepts both positive and negative prompts generated by DPG. With the CFG strategy, the negative prompt replaces the unconditional input and guides the model to learn how to approach HQ features while avoiding LQ characteristics. In addition, to further evaluate performance under both generic and HMB degradations, we construct a new benchmark, MPII-Test. It is curated from the MPII Human Pose dataset andriluka2014mpii and contains 5,427 real-world degraded human images, many of which include rich motion blur patterns. As shown in Fig. 1, the results demonstrate our model’s strong ability to restore human images with both HMB and other degradations.
@@ 2.1 Human Body Restoration
Human body restoration (HBR) constitutes a specialized subfield that applies blind image restoration (BIR) techniques specifically to the restoration of LQ human images. Contemporary mainstream BIR methods liang2021swinir ; wang2021real ; zhang2021bsrgan ; lin2024diffbir show remarkable efficacy across a wide range of natural scenarios. However, their direct application to human body images frequently causes joint misalignment and limb distortion. Previous models Liu2021HumanBodySR ; Wang2024PRCN ; Zhang2024DiffBody ; gong2025osdhuman have addressed this task with varying methodologies. Among them, DiffBody Zhang2024DiffBody pioneers the application of diffusion model in body-region image enhancement under the guidance of the body attention module. A notable contribution to the field is OSDHuman gong2025osdhuman , which proposes a one-step diffusion model for human body restoration using a high-fidelity image embedder while introducing the PERSONA benchmark dataset.
Human motion blur represents one of the most challenging degradation modalities in HBR contexts. Current deep-learning-based deblurring research primarily focuses on objects and scenes  nah2017deep ; chen2023hierarchical ; tao2018scale ; Kupyn2018DeblurGAN . DeblurGAN Kupyn2018DeblurGAN represents the inaugural application of a conditional adversarial network to blind motion deblurring and introduced a random-trajectory simulation pipeline for synthesizing motion-blurred data. More recently, OSDD liu2025OSDD , a one-step diffusion model for motion deblurring that substantially enhances the computational efficiency of diffusion-based restoration. Moreover, ReLoBlur li2023real provides the first real-world locally blurred dataset captured with synchronized light-field cameras, while developing a Blur-Aware Gating network to restore these regions.
@@ 2.2 Diffusion Models
Text-to-image (T2I) diffusion models are repurposed for image restoration tasks due to their powerful prior knowledge in image generation wang2024exploiting ; lin2024diffbir ; wu2024seesr ; yu2024supir ; yang2023pasd . For instance, Stable Diffusion (SD) sd21 , with scalable networks and controllable generation, demonstrates the capability to inject vivid details into LQ images. Building on this foundation, StableSR wang2024exploiting enhances image restoration via a fine-tuned time-aware encoder and progressive sampling strategies. By implementing a degradation-aware prompt extractor, SeeSR wu2024seesr guides diffusion models to generate semantically accurate HQ images with precise prompt control. DiffBIR xia2023diffir first employs an initial restoration module before incorporating SD for detail refinement. Furthermore, SUPIR yu2024supir leverages SDXL podell2023sdxl as its prior, achieving impressive results through high-quality datasets and innovative positive-negative sample strategies. However, their adherence to conventional T2I diffusion paradigms necessitates numerous sampling steps, resulting in computational inefficiency and excessive parametric complexity.
Contemporary research increasingly focuses on addressing the inefficiencies of multi-step diffusion processes in BIR tasks by leveraging more efficient one-step diffusion wang2024sinsr ; wu2024osediff ; 2024s3diff ; yue2024invsr . SinSR wang2024sinsr advances this direction by distilling a deterministic mapping from a teacher diffusion model. OSEDiff wu2024osediff injects LQ images into the latent space as the diffusion starting point and employs variational score distillation to align with the image prior of SD. More recently, based on diffusion inversion, InvSR yue2024invsr uses noise prediction to create optimal intermediate sampling states. Nevertheless, they insufficiently exploit the semantic information inherent within the LQ images themselves. The global semantic content of LQ images can be effectively processed and utilized as prompt embeddings for diffusion models, substantially enhancing one-step diffusion restoration capabilities.
@@ 3.1 Degradation Pipeline with Human Motion Blur (HMB)
Obtaining paired training data where degradation incorporates HMB remains a significant challenge. Prior works nah2017deep ; noroozi2017motion ; li2023real construct global or local motion blur by capturing high-quality continuous frames from high frame-rate videos and averaging them to synthesize motion blur. However, this strategy struggles to cover the diversity of real-world human activities, limiting the generalizability of the dataset. Another method involves convolving natural images with blur kernels generated from complex motion trajectories chakrabarti2016neural ; sun2015learning ; xu2014deep ; boracchi2012modeling , as exemplified by DeblurGAN Kupyn2018DeblurGAN . Building upon these methods, we design a degradation pipeline that simulates human motion blur and incorporates generic degradation processes based on Real-ESRGAN wang2021real to train for human body restoration.
@@
The pipeline is illustrated in Fig. 2. In the first order of degradation, we use Sapiens khirodkar2024sapiens for body-part segmentation on HQ images $I_{H}$, yielding part masks. These masks are grouped into six categories (head, left/right upper limbs, left/right lower limbs, and the whole body), from which one category is randomly selected for motion blur simulation in subsequent steps. Since segmented regions do not inherently correspond to realistic motion patterns, we apply morphological operations, including erosion, dilation, and Gaussian blurring, to the selected masks. These processed masks are then normalized to generate a spatial weight map $W_{\text{s}}$, which can be formalized as:
[…]
@@
In parallel, we generate a globally blurred image using a strategy similar to DeblurGAN Kupyn2018DeblurGAN : a continuous random trajectory is simulated via a Markov process and converted to a discrete point spread function (PSF) through bilinear interpolation. The PSF is then convolved with the HQ image via FFT (fast Fourier transform, denoted as $\mathcal{F}$) convolution to produce global motion blur. Finally, we blend the original and blurred images using the spatial weight map, yielding the HMB image $I_{\text{HMB}}$:
[…]
@@
Notably, HMB is typically caused by the subject’s movement during capture and should logically precede all other degradations in the simulation process. Therefore, to maintain logical consistency, we place HMB in the first-order stage. Since motion blur and severe degradations do not always occur, we define the first-order degradation as one of three possible conditions: no degradation, HMB, or generic degradation. After this stage, the image undergoes second-order degradation, which includes common types such as blur, resizing, noise, and compression. The generic degradations in both the first and second orders follow the strategy of Real-ESRGAN wang2021real . Upon traversing this comprehensive degradation pipeline, we obtain synthesized LQ images $I_{L}$ for subsequent processing.
@@
A key challenge in prompt extraction is effectively predicting high-quality (HQ) features from low-quality (LQ) images. OSDFace wang2024osdface tackles this via HQ–LQ embedding alignment. However, alignment alone is insufficient because predicting HQ features from LQ involves restoration, which is challenging for prompt extractors. Directly training the prompt extractor with LQ inputs and HQ targets embeds HQ features within intermediate model layers. Additionally, for restoration models, defining the negative prompt is essential: the restored images should not completely diverge from the LQ input but rather specifically from the residual noise (LQ minus HQ), which contains degradation without structural information. Otherwise, the restored images may lack fidelity. Furthermore, local noise (e.g., human motion blur) hidden in global noise is hard to locate and remove. Existing methods li2023real ; li2024adaptive use segmentation supervision to locate motion blur regions. A dedicated module that supplies explicit positional information can further enhance noise localization and removal.
DPG Structure. To provide the model with image embeddings as prompts, the Vision Transformer dosovitskiy2020vit ; liu2021Swin framework is suitable. It can convert RGB images into embeddings for feature extraction. Furthermore, SwinIR liang2021swinir in image restoration, using composed residual Swin Transformer blocks (RSTB) to recover HQ from LQ, achieves great success. Inspired by SwinIR, our prompt extractor structure is shown in Fig. 3. First, we use a convolutional structure ($H_{D}$). It downsamples the LQ image ($I_{L}$) size 512$\times$512 by 4. And it increases channels to match the sequence length by the Swin Transformer, which we set to 150. Next, the backbone network (denoted $H_{E}$) extracts features through two RSTBs, with six Swin Transformer layers (STL). Afterward, the model splits into three reconstruction branches (denoted as $H_{Ri},i\in[1,3]$). Each branch also contains two RSTBs but only three STLs. Each branch then passes through a convolutional upsampler (denoted as $H_{Ui},i\in[1,3]$). The three branches are designed to predict the HQ image ($\hat{I}_{H}^{P}$), the residual noise between LQ and HQ ($\hat{I}_{R}$), and the HMB segmentation mask ($\hat{M}_{\text{HMB}}$). This process can be formulated as follows:
@@
Training Objective of DPG. The predicted HQ image ($\hat{I}_{H}^{P}$) and the residual noise between LQ and HQ ($\hat{I}_{R}$) are optimized jointly using the pixel-wise L1 loss $\mathcal{L}_{1}$. The HMB-aware branch, responsible for predicting the human motion blur (HMB) segmentation mask ($\hat{M}_{\text{HMB}}$), uses the Dice dice1945ecologic loss $\mathcal{L}_{\text{Dice}}$. The overall training objective can be formulated as follows:
@@
Model Structure. The latent diffusion model (LDM) Rombach2022LDM adds noise in latent space at timestep $t$ as $z_{t}=\sqrt{\bar{\alpha}_{t}}\,z+\sqrt{1-\bar{\alpha}_{t}}\,\varepsilon$, where $\varepsilon\sim\mathcal{N}(0,I)$ and $\bar{\alpha}_{t}$ denotes the cumulative product of ${\alpha}_{s}$ up to timestep $t$: $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$. The reverse process predicts noise with parameter $\theta$. During inference, the LDM employs the DDIM song2021ddim to accelerate the reverse process, simplified as follows:
@@
where $t-1$ represents the next step in the DDIM sampling sequence. The $\sigma_{t}^{2}$ is defined as follows:
@@
In restoration tasks, for stability, we set $\eta{=}0$, so $\sigma_{t}{=}0$. With a specific timestep $\tau$, the latent vector $z_{\tau}$ should correspond to the noisy representation $z_{L}$ encoded from the LQ image $I_{L}$ by the variational autoencoder (VAE) kingma2014vae encoder $E_{\theta}$. $t-1$ should correspond to step 0 in OSD. Thus $\bar{\alpha}_{t-1}=1$, since the output is the noise-free latent vector $z_{0}$, also denoted as $\hat{z}_{H}$: $\hat{z}_{H}=(z_{t}-\sqrt{1-\bar{\alpha}_{t}}\,z_{\varepsilon})/{\sqrt{\bar{\alpha}_{t}}}$. When the classifier-free guidance (CFG) ho2021cfg is employed, the predicted noise $z_{\varepsilon}$ is defined as:

### v2

@@ Abstract
Human-centered images often suffer from severe generic degradation during transmission and are prone to human motion blur (HMB), making restoration challenging. Existing research lacks sufficient focus on these issues, as both problems often coexist in practice. To address this, we design a degradation pipeline that simulates the coexistence of HMB and generic noise, generating synthetic degraded data to train our proposed HAODiff, a human-aware one-step diffusion. Specifically, we propose a triple-branch dual-prompt guidance (DPG), which leverages high-quality images, residual noise (LQ minus HQ), and HMB segmentation masks as training targets. It produces a positive–negative prompt pair for classifier‑free guidance (CFG) in a single diffusion step. The resulting adaptive dual prompts let HAODiff exploit CFG more effectively, boosting robustness against diverse degradations. For fair evaluation, we introduce MPII‑Test, a benchmark rich in combined noise and HMB cases. Extensive experiments show that our HAODiff surpasses existing state-of-the-art (SOTA) methods in terms of both quantitative metrics and visual quality on synthetic and real-world datasets, including our introduced MPII-Test. Code is available at: https://github.com/gobunu/HAODiff.
@@ 1 Introduction
Human body restoration (HBR) focuses on recovering high-quality (HQ) images from low-quality (LQ) inputs that contain human images. When human subjects appear prominently in an image, it naturally attracts more viewer attention. However, real-world images frequently undergo degradation during capture or transmission, including human motion blur, noise, resolution loss, and JPEG compression artifacts. These distortions significantly hinder the recognition of human activities and limit the usefulness of such images in broader applications. As a result, many downstream tasks related to humans are negatively affected, such as 3D reconstruction Wang et al. (2021a); Zhang et al. (2025), human pose estimation Samkari et al. (2023); Zheng et al. (2025), and human-object interaction detection Wang et al. (2024d); Liu et al. (2025b).
To achieve better performance in practical scenarios, current blind image restoration (BIR) models typically rely on a large number of paired LQ and HQ images to accurately learn the complex mapping between LQ and HQ domains. Models in the HBR field follow the same principle. However, collecting a large number of real LQ-HQ image pairs is challenging. Real-world LQ images often undergo various unknown degradations, which may occur during transmission or even at the time of capture. Therefore, BIR models usually adopt a degradation pipeline to simulate real-world LQ conditions. This strategy allows models to fully leverage large-scale high-quality image restoration datasets, such as LSDIR Li et al. (2023b), FFHQ Karras et al. (2019), and PERSONA Gong et al. (2025). In the context of HBR, existing approaches Gong et al. (2025); Zhang et al. (2024b) generally use generic degradation types, including downsampling, compression, noise, and low-pass blur. A commonly adopted pipeline is from Real-ESRGAN Wang et al. (2021b), which is originally developed for blind super-resolution tasks. However, this pipeline may fail to cover the diverse degradation types that frequently occur in human-centric images.
@@
Among the various degradation types in human-centric images, human motion blur is particularly prevalent, yet it remains significantly underrepresented in existing degradation pipelines. Motion blur can be broadly categorized into two types. The first category is global motion blur, typically caused by camera movement during exposure. The second is local motion blur, caused by the rapid movement of objects in the scene. In human images, local motion blur mainly arises from human movement, known as human motion blur (HMB). Extensive research is conducted on the removal of global motion blur. For example, DeblurGAN Kupyn et al. (2018) not only proposes an effective model for motion deblurring, but also provides a simulation pipeline that can generate realistic global motion blur. This enables training with large-scale synthetic data. In contrast, the restoration of local motion blur is typically based on real datasets, with ReLoBlur Li et al. (2023a) being a representative example that covers diverse object motion. However, these datasets often focus on general objects and pay less attention to human-specific motion blur. Moreover, HMB in real-world human images often co-occurs with other degradation types. Existing deblurring models are usually designed to handle a single degradation, which limits their applicability in more complex, realistic scenarios.
@@
Multiple factors, including training data and degradation pipelines, influence the quality of image restoration. Among them, the model architecture plays a critical role. In BIR tasks, models require strong generative capabilities to reconstruct damaged regions. Generative adversarial networks (GANs) Goodfellow et al. (2014) are considered one of the major starting points for modern image generation models. They have led to the development of many powerful restoration methods Zhang et al. (2021); Kupyn et al. (2018); Wang et al. (2021b). However, GANs often suffer from training instability and difficulties in controlling the generation process. Diffusion models and latent diffusion models (LDMs) Rombach et al. (2022) have further improved the framework for both generation and restoration. Some models Lin et al. (2024); Yang et al. (2023); Wu et al. (2024b); Yue et al. (2023) use multi-step diffusion processes to restore high-quality images from heavily degraded inputs. Recently, one-step diffusion models Wu et al. (2024a); Wang et al. (2025, 2024c); Yue et al. (2025) also demonstrate strong BIR performance while significantly reducing resource consumption compared to multi-step diffusion. These models commonly extract object features from images and convert them into text or image embeddings. They are used as positive prompts to guide the model toward faithful restoration. To better leverage the guidance capabilities of text-to-image (T2I) foundation models, some methods Zhang et al. (2024a); Yu et al. (2024) introduce negative prompts using classifier-free guidance (CFG) Ho and Salimans (2021). These prompts help steer the model away from undesired content. However, these methods adopt fixed negative prompts, which limit their guidance effectiveness.
@@
To address these limitations, we propose HAODiff, a novel one-step diffusion for human body restoration (HBR) that integrates CFG via dual-prompt guidance (DPG). Firstly, to compensate for the lack of human motion blur (HMB) in existing pipelines, we propose a new one. In the preprocessing stage, we apply a human body-part segmentation model to the HQ training images to obtain part-specific masks. These masks are then combined with a motion blur simulation module to synthesize HMB. Integrated into a generic two‑stage pipeline, our approach allows the inclusion of HMB. Secondly, we design a triple-branch dual-prompt guidance, named DPG, which is based on the Swin Transformer Liu et al. (2021b). During training with the degradation pipeline, one branch of DPG predicts the HQ image to produce the positive prompt. The other two are used to predict the residual noise (LQ minus HQ) and the human motion blur segmentation mask, both serving as sources for the negative prompt. Thirdly, we propose a human-aware one-step diffusion via dual-prompt guidance, named HAODiff. The model accepts both positive and negative prompts generated by DPG. With the CFG strategy, the negative prompt replaces the unconditional input and guides the model to learn how to approach HQ features while avoiding LQ characteristics. In addition, to further evaluate performance under both generic and HMB degradations, we construct a new benchmark, MPII-Test. It is curated from the MPII Human Pose dataset Andriluka et al. (2014) and contains 5,427 real-world degraded human images, many of which include rich motion blur patterns. As shown in Fig. 1, the results demonstrate our model’s strong ability to restore human images with both HMB and other degradations.
@@ 2.1 Human Body Restoration
Human body restoration (HBR) constitutes a specialized subfield that applies blind image restoration (BIR) techniques specifically to the restoration of LQ human images. Contemporary mainstream BIR methods Liang et al. (2021); Wang et al. (2021b); Zhang et al. (2021); Lin et al. (2024) show remarkable efficacy across a wide range of natural scenarios. However, their direct application to human body images frequently causes joint misalignment and limb distortion. Previous models Liu et al. (2021a); Wang et al. (2024b); Zhang et al. (2024b); Gong et al. (2025) have addressed this task with varying methodologies. Among them, DiffBody Zhang et al. (2024b) pioneers the application of diffusion model in body-region image enhancement under the guidance of the body attention module. A notable contribution to the field is OSDHuman Gong et al. (2025), which proposes a one-step diffusion model for human body restoration using a high-fidelity image embedder while introducing the PERSONA benchmark dataset.
Human motion blur represents one of the most challenging degradation modalities in HBR contexts. Current deep-learning-based deblurring research primarily focuses on objects and scenes  Nah et al. (2017); Chen et al. (2023); Tao et al. (2018); Kupyn et al. (2018); Cao et al. (2025). DeblurGAN Kupyn et al. (2018) represents the inaugural application of a conditional adversarial network to blind motion deblurring and introduced a random‑trajectory simulation pipeline for synthesizing motion‑blurred data. More recently, OSDD Liu et al. (2025a), a one‑step diffusion model for motion deblurring that substantially enhances the computational efficiency of diffusion‑based restoration. Moreover, ReLoBlur Li et al. (2023a) provides the first real-world locally blurred dataset captured with synchronized light-field cameras, while developing a Blur-Aware Gating network to restore these regions.
@@ 2.2 Diffusion Models
Text-to-image (T2I) diffusion models are repurposed for image restoration tasks due to their powerful prior knowledge in image generation Wang et al. (2024a); Lin et al. (2024); Wu et al. (2024b); Yu et al. (2024); Yang et al. (2023). For instance, Stable Diffusion (SD) Stability AI (2022), with scalable networks and controllable generation, demonstrates the capability to inject vivid details into LQ images. Building on this foundation, StableSR Wang et al. (2024a) enhances image restoration via a fine-tuned time-aware encoder and progressive sampling strategies. By implementing a degradation-aware prompt extractor, SeeSR Wu et al. (2024b) guides diffusion models to generate semantically accurate HQ images with precise prompt control. DiffBIR Xia et al. (2023) first employs an initial restoration module before incorporating SD for detail refinement. Furthermore, SUPIR Yu et al. (2024) leverages SDXL Podell et al. (2023) as its prior, achieving impressive results through high-quality datasets and innovative positive-negative sample strategies. However, their adherence to conventional T2I diffusion paradigms necessitates numerous sampling steps, resulting in computational inefficiency and excessive parametric complexity.
Contemporary research increasingly focuses on addressing the inefficiencies of multi-step diffusion processes in BIR tasks by leveraging more efficient one-step diffusion Wang et al. (2024c); Wu et al. (2024a); Zhang et al. (2024a); Yue et al. (2025). SinSR Wang et al. (2024c) advances this direction by distilling a deterministic mapping from a teacher diffusion model. OSEDiff Wu et al. (2024a) injects LQ images into the latent space as the diffusion starting point and employs variational score distillation to align with the image prior of SD. More recently, based on diffusion inversion, InvSR Yue et al. (2025) uses noise prediction to create optimal intermediate sampling states. Nevertheless, they insufficiently exploit the semantic information inherent within the LQ images themselves. The global semantic content of LQ images can be effectively processed and utilized as prompt embeddings for diffusion models, substantially enhancing one-step diffusion restoration capabilities.
@@ 3.1 Degradation Pipeline with Human Motion Blur (HMB)
Obtaining paired training data where degradation incorporates HMB remains a significant challenge. Prior works Nah et al. (2017); Noroozi et al. (2017); Li et al. (2023a) construct global or local motion blur by capturing high-quality continuous frames from high frame-rate videos and averaging them to synthesize motion blur. However, this strategy struggles to cover the diversity of real-world human activities, limiting the generalizability of the dataset. Another method involves convolving natural images with blur kernels generated from complex motion trajectories Chakrabarti (2016); Sun et al. (2015); Xu et al. (2014); Boracchi and Foi (2012), as exemplified by DeblurGAN Kupyn et al. (2018). Building upon these methods, we design a degradation pipeline that simulates human motion blur and incorporates generic degradation processes based on Real-ESRGAN Wang et al. (2021b) to train for human body restoration.
@@
The pipeline is illustrated in Fig. 2. In the first order of degradation, we use Sapiens Khirodkar et al. (2024) for body‑part segmentation on HQ images $I_{H}$, yielding part masks. These masks are grouped into six categories (head, left/right upper limbs, left/right lower limbs, and the whole body), from which one category is randomly selected for motion blur simulation in subsequent steps. Since segmented regions do not inherently correspond to realistic motion patterns, we apply morphological operations, including erosion, dilation, and Gaussian blurring, to the selected masks. These processed masks are then normalized to generate a spatial weight map $W_{\text{s}}$, which can be formalized as:
$W_{\text{s}}=(\mathit{Norm}\circ\mathit{Morph}\circ\mathit{Seg})(I_{H})\quad\text{with}\quad\mathit{Morph}=G\circ D\circ E,$ (1)
where $\mathit{Seg}(\cdot)$ refers to the segmentation operation, and $\mathit{Morph}(\cdot)$ denotes morphological operations: Gaussian blur ($G$), dilation ($D$) and erosion ($E$). $\mathit{Norm}(\cdot)$ denotes scaling values to $[0,1]$.
@@
In parallel, we generate a globally blurred image using a strategy similar to DeblurGAN Kupyn et al. (2018): a continuous random trajectory is simulated via a Markov process and converted to a discrete point spread function (PSF) through bilinear interpolation. The PSF is then convolved with the HQ image via FFT (fast Fourier transform, denoted as $\mathcal{F}$) convolution to produce global motion blur. Finally, we blend the original and blurred images using the spatial weight map, yielding the HMB image $I_{\text{HMB}}$:
$I_{\text{HMB}}=W_{s}\odot I_{H}+(1-W_{s})\odot I_{B}\quad\text{with}\quad I_{B}=\mathcal{F}^{-1}\left(\mathcal{F}(\mathit{PSF})\odot\mathcal{F}(I_{H})\right).$ (2)
@@
Notably, HMB is typically caused by the subject’s movement during capture and should logically precede all other degradations in the simulation process. Therefore, to maintain logical consistency, we place HMB in the first-order stage. Since motion blur and severe degradations do not always occur, we define the first-order degradation as one of three possible conditions: no degradation, HMB, or generic degradation. After this stage, the image undergoes second-order degradation, which includes common types such as blur, resizing, noise, and compression. The generic degradations in both the first and second orders follow the strategy of Real-ESRGAN Wang et al. (2021b). Upon traversing this comprehensive degradation pipeline, we obtain synthesized LQ images $I_{L}$ for subsequent processing.
@@
A key challenge in prompt extraction is effectively predicting high-quality (HQ) features from low-quality (LQ) images. OSDFace Wang et al. (2025) tackles this via HQ–LQ embedding alignment. However, alignment alone is insufficient because predicting HQ features from LQ involves restoration, which is challenging for prompt extractors. Directly training the prompt extractor with LQ inputs and HQ targets embeds HQ features within intermediate model layers. Additionally, for restoration models, defining the negative prompt is essential: the restored images should not completely diverge from the LQ input but rather specifically from the residual noise (LQ minus HQ), which contains degradation without structural information. Otherwise, the restored images may lack fidelity. Furthermore, local noise (e.g., human motion blur) hidden in global noise is hard to locate and remove. Existing methods Li et al. (2023a, 2024) use segmentation supervision to locate motion blur regions. A dedicated module that supplies explicit positional information can further enhance noise localization and removal.
DPG Structure. To provide the model with image embeddings as prompts, the Vision Transformer Dosovitskiy et al. (2021); Liu et al. (2021b) framework is suitable. It can convert RGB images into embeddings for feature extraction. Furthermore, SwinIR Liang et al. (2021) in image restoration, using composed residual Swin Transformer blocks (RSTB) to recover HQ from LQ, achieves great success. Inspired by SwinIR, our prompt extractor structure is shown in Fig. 3. First, we use a convolutional structure ($H_{D}$). It downsamples the LQ image ($I_{L}$) size $\times512$ by 4. And it increases channels to match the sequence length by the Swin Transformer, which we set to 150. Next, the backbone network (denoted $H_{E}$) extracts features through two RSTBs, with six Swin Transformer layers (STL). Afterward, the model splits into three reconstruction branches (denoted as $H_{Ri},i\in[1,3]$). Each branch also contains two RSTBs but only three STLs. Each branch then passes through a convolutional upsampler (denoted as $H_{Ui},i\in[1,3]$). The three branches are designed to predict the HQ image ($\hat{I}_{H}^{P}$), the residual noise between LQ and HQ ($\hat{I}_{R}$), and the HMB segmentation mask ($\hat{M}_{\text{HMB}}$). This process can be formulated as follows:
@@
Training Objective of DPG. The predicted HQ image ($\hat{I}_{H}^{P}$) and the residual noise between LQ and HQ ($\hat{I}_{R}$) are optimized jointly using the pixel‑wise L1 loss $\mathcal{L}_{1}$. The HMB-aware branch, responsible for predicting the human motion blur (HMB) segmentation mask ($\hat{M}_{\text{HMB}}$), uses the Dice Dice (1945) loss $\mathcal{L}_{\text{Dice}}$. The overall training objective can be formulated as follows:
@@
Model Structure. The latent diffusion model (LDM) Rombach et al. (2022) adds noise in latent space at timestep $t$ as $z_{t}=\sqrt{\bar{\alpha}_{t}}\,z+\sqrt{1-\bar{\alpha}_{t}}\,\varepsilon$, where $\varepsilon\sim\mathcal{N}(0,I)$ and $\bar{\alpha}_{t}$ denotes the cumulative product of ${\alpha}_{s}$ up to timestep $t$: $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$. The reverse process predicts noise with parameter $\theta$. During inference, the LDM employs the DDIM Song et al. (2021) to accelerate the reverse process, simplified as follows:
@@
where t1t-1 represents the next step in the DDIM sampling sequence. The $\sigma_{t}^{2}$ is defined as follows:
@@
In restoration tasks, for stability, we set $\eta{=}0$, so $\sigma_{t}{=}0$. With a specific timestep $\tau$, the latent vector $z_{\tau}$ should correspond to the noisy representation $z_{L}$ encoded from the LQ image $I_{L}$ by the variational autoencoder (VAE) Kingma and Welling (2014) encoder $E_{\theta}$. t1t-1 should correspond to step 0 in OSD. Thus $\bar{\alpha}_{t-1}=1$, since the output is the noise-free latent vector $z_{0}$, also denoted as $\hat{z}_{H}$: $\hat{z}_{H}=(z_{t}-\sqrt{1-\bar{\alpha}_{t}}\,z_{\varepsilon})/{\sqrt{\bar{\alpha}_{t}}}$. When the classifier-free guidance (CFG) Ho and Salimans (2021) is employed, the predicted noise $z_{\varepsilon}$ is defined as: